please help if you can,
Solve this system of equations
4a+7b=27
8a-3c=-22
6b-3c=24
The solution is [],[],[] Thank you

Answer by scott8148(6628)

A) third equation minus second equation ... (6b-3c)-(8a-3c)=24-(-22) ... 6b-8a=46

B) two times second equation ... 2(4a+7b=27) ... 8a+14b=54

C) A+B ... (6b-8a)+(8a+14b)=46+54 ... 20b=100 ... b=5

from second equation ... 4a+(7*5)=27 ... 4a=-8 ... a=-2

from first equation ... (6*5)-3c=24 ... -3c=-6 ... c=2